A 4.98 g sample of aniline (C6H5NH2, molar mass = 93.13 g/mol) was combusted in a bomb

calorimeter with a heat capacity of 4.25 kJ/
Answer #1

Heat released by combustion raises temerature in calorimeter

heat released=heat capacity*rise in temperature

         = 4.25*(69.8-29.5)

        =171.275 KJ

mol of aniline=mass/moloar mass

                =4.98/93.13

                =0.0535 mol

Delta H=-heat released/mol

            =-171.275/0.0535

           =-3201.4 Kj/mol